### v2

@@ Abstract
This paper provides a detailed examination of the advancements and applications of large language models in the healthcare sector, with a particular emphasis on clinical applications. The study traces the evolution of LLMs from their foundational technologies to the latest developments in domain-specific models and multimodal integration. It explores the technical progression from encoder-based models requiring fine-tuning to sophisticated approaches that integrate textual, visual, and auditory data, thereby facilitating comprehensive AI solutions in healthcare. The paper discusses both the opportunities these technologies present for enhancing clinical efficiency and the challenges they pose in terms of ethics, data privacy, and implementation. Additionally, it critically evaluates the deployment strategies of LLMs, emphasizing the necessity of open-source models to ensure data privacy and adaptability within healthcare environments. Future research directions are proposed, focusing on empirical studies to evaluate the real-world efficacy of LLMs in healthcare and the development of open datasets for further research. This review aims to provide a comprehensive resource for both newcomers and multidisciplinary researchers interested in the intersection of AI and healthcare.
Introduction
@@
The development of artificial intelligence (AI) in recent years has opened countless opportunities for various sectors, including healthcare. The potential influence of AI is a subject of debate concerning its impact on humanity. Leading AI experts have called for caution, evidenced by an open letter urging a pause in the expansion of advanced AI models, which reflects growing concerns among policymakers and the public about the ethical, social, and economic ramifications of AI. While some argue that AI can bring substantial advances in efficiency and effectiveness across many sectors, others fear it could exacerbate inequalities, displace jobs, and challenge societal norms [1]. While AI in healthcare has an extensive history of research [2], the emergence of advanced language and multimodal models such as GPT family[3], Gemini [4], and a series of open models like the Llama [5], offers unprecedented perspectives for the transformation of the healthcare sector.
This review paper synthesizes and critically examines the landscape of large language models (LLMs) within the medical domain, focusing on their clinical applications. Many existing survey papers cover LLMs in medicine. The paper titled “Pre-trained Language Models in Medicine: A Survey” [6] offers a broad overview of applications but may prove challenging for newcomers due to the wide range of topics encompassed. Similarly, “Pre-trained Language Models in Biomedical Domain: A Systematic Survey” [7] delves deeply into the textual applications of LLMs in biomedicine. However, its extensive coverage, which encompasses applications beyond the clinical subdomain, and the complex structure of the article limit its applicability for novice or multidisciplinary researchers. Finally, “Large Language Models in Medicine: Current Potential and Opportunities for Development” [8] highlights the capabilities of advanced textual models such as GPT-3.5 and GPT-4. Although the review comprehensively describes the potential applications of LLMs, categorizing them into clinical, research, and educational purposes, it lacks important implementation details and a detailed exploration of open-source models.
The purpose of this review is to systematically present the capabilities of LLMs in the medical domain, with a primary focus on clinical applications. This review deliberately avoids extensive discussions on their use in medical education and research. Instead, it aims to integrate these models with other approaches and identify prevailing trends, with an emphasis on textual models. The goal is to clearly outline the practical applications of these technologies, making them accessible and actionable for both researchers and practitioners.
@@ 1.1 Language Models
A critical milestone in NLP was the introduction of the attention mechanism in neural machine translation [9]. This mechanism, linking the encoder and decoder in sequence-to-sequence models, paved the way for subsequent advancements by enabling the model to focus on different parts of the input sequence for each step of the output, substantially improving the handling of longer input sequences and complex dependencies. Notably, it led to the creation of the Transformer model [10], which exclusively relies on attention mechanisms. This innovation revolutionized not only the field of NLP but also the broader realms of AI and machine learning (ML).
The emergence of Transformer-based models has spurred the development of a wide array of large language models, both in commercial and open-source formats. The most prominent series of these models, GPT [11], is based on the decoder part of the original Transformer architecture. Decoder-based models are particularly suitable for autoregressive text generation. In contrast, models like BERT [12] and its variants, which are encoder-based, are typically trained using a masked language modeling approach. While they are arguably less suitable for text generation, their applications are vast as they can be customized and fine-tuned for a plethora of tasks, including question answering, sentiment analysis, information extraction, and many more. There are also models that integrate both encoder and decoder components, such as T5 [13] and BART [14]. These models can effectively transform input sequences into output sequences, thus naturally fitting tasks such as summarization and paraphrasing.
@@
Despite the simplicity of statistical approaches to language modeling and their initial lack of focus on the underlying rules of language, as critiqued by Chomsky [15], contemporary language models like GPT-4 have exhibited remarkable proficiency in various language understanding tasks [3].
These models demonstrate emergent abilities and adeptness in learning natural language patterns. Modern language models have undergone a notable paradigm shift, moving from a pre-training and fine-tuning approach to embracing in-context learning (ICL) [16] and zero-shot learning. This advancement enables them to solve tasks without the need for updating the models’ weights. Nevertheless, challenges remain, as underscored by Mahowald et al. [17]. They highlight the gap between formal and functional linguistic competencies in LLMs, noting these models often lack functional competencies. They propose integrating modularity into the system architecture or attaining it through a revised training process.
Notably, current research is shifting towards multimodal models that can integrate multiple modalities: visual, text, and potentially audio, into a single model [18], thus paving the way for more comprehensive AI solutions.
1.2 Open-source LLMs
Open-source LLMs are publicly accessible and, unlike their commercial counterparts, are freely available for modification and distribution. This openness allows for customization and adaptation to specific needs, including deployment on private servers, which enhances data privacy and ownership, paramount in sensitive fields like medicine.
In the medical domain, the ability to deploy LLMs on-premise is a substantial advantage. It ensures that sensitive medical data does not have to be shared with third-party providers, thereby complying with stringent data protection regulations. Healthcare organizations can also tailor these models to better understand and generate medical language, optimize them for specific types of medical inquiries, and integrate them with confidential datasets, all while maintaining patient confidentiality.
@@ 1.2 Open-source LLMs
However, open-source LLMs have limitations compared to commercial models. They often lack the extensive support, continuous updates, and robust infrastructure provided by commercial vendors. These deficiencies can lead to challenges in model maintenance, scalability, and performance. The primary barrier to the widespread adoption of advanced open-source LLMs is their significant number of parameters, each typically represented by a floating-point number. Generally, a higher number of parameters enhances reasoning, pattern recognition, and linguistic capabilities, but this comes at the cost of increased storage requirements and the need for more powerful GPUs, both for inference and fine-tuning. Table 1 summarizes widely used pre-trained open-source textual and multimodal models (predominantly English-language) categorized by the number of parameters and their architectures.
Recent research has provided notable solutions for mitigating the computational resource needs of large models by introducing parameter-efficient fine-tuning techniques and reducing memory and storage requirements. For instance, Low-Rank Adaptation (LoRA) [50] optimizes updates through low-rank parameter adjustments instead of updating all model parameters. Additionally, applying quantization techniques [51] allows open-source LLMs to reduce the memory footprint and enhance inference speed, thus making them more efficient and suitable for deployment across a broader range of hardware. Quantized Low-Rank Adaptation (QLoRA) [52] combines these techniques to efficiently fine-tune large language models while significantly reducing memory and computational demands. For example, a model with 1 billion parameters, when using 16-bit floating-point precision, requires at least 2GB of GPU memory for inference, depending on the input length. In contrast, a quantized version of the same model could require approximately half of this amount. Furthermore, recent work by Guo et al. proposes a novel memory allocation framework that utilizes low-level GPU virtual memory management. This framework incorporates a Virtual Memory Stitching (VMS) mechanism, reducing GPU memory usage and fragmentation, thereby further enhancing the efficiency of deploying LLMs [53]. Another notable advancement, flash attention [54], allows for more efficient hardware usage thus opening additional opportunities for on-premise hardware to utilize larger models.
@@ 1.3 Domain-specific Language Models in Healthcare
Domain-specific LLMs are models tailored for a narrow field or topic, enhancing their ability to produce more accurate and relevant responses within that specific area. Generally, there are two common methods to create a domain-specific model: one approach is to pre-train a model using a set of domain-specific documents, such as medical papers; another is to take a generically trained model and fine-tune or adapt it to the target domain.
The medical field, rich in unstructured textual data from sources such as Electronic Health Records (EHR) and other medical documentation, presents an ideal research domain for language models to address a wide array of challenges. These applications range from tasks previously tackled by NLP techniques, such as clinical acronym disambiguation, to those unattainable a decade ago. For example, medical chatbots that can assist both patients and healthcare professionals are now emerging. In response to these advancements, the research community has been actively developing language models specifically designed for medical applications. The evolution of these models, from pre-training and fine-tuning strategies to creating open-source models capable of zero-shot learning, highlights the growing synergy between artificial intelligence and healthcare [88, 94, 113, 95].
@@
The debut of BioGPT has marked another milestone in biomedical NLP. This generative pre-trained Transformer, based on GPT-2 and trained on millions of PubMed abstracts, has demonstrated superior performance in specialized tasks, including text classification, biomedical text generation, and data mining [85].
@@
AlpaCare, a medical instruction-tuned LLM trained on more than fifty thousand instruction-response pairs generated artificially using GPT-4, represents a notable example from the current generation of medical LLMs that can be utilized in the medical domain without the need for fine-tuning on downstream tasks [88]. Clinical Camel [95], another medical LLM fine-tuned from the Llama-2 model, excels in various medical tasks, ranging from clinical note creation to medical triaging. Despite its limitations, such as the potential for generating misleading content and the need for continuous updates, it marked a remarkable advancement in medical LLMs. The issue of needing continuous updates was addressed by the ChatDoctor [94] model. Based on the Llama family and fine-tuned with real-world patient-doctor conversations, this model can access external information sources, a crucial capability in the medical field, especially for emerging diseases.
@@
Recent advancements in the field, particularly the development of multimodal models that work with both text and images, have substantially impacted the medical domain as well. The LLaVA-Med model [101] can effectively answer open-ended questions about biomedical images, while the CheXagent model [105], designed specifically for chest X-ray interpretation, exemplifies the successful transition from general medical models to specialist models capable of effectively tackling problems in narrow medical fields.
@@
Table 2 provides a comprehensive, though not exhaustive, summary of existing medical LLMs. It illustrates current trends in the development of medical LLMs, as well as the evolution of training and testing datasets utilized.
2 Medical Applications of LLMs

### v3

@@ Abstract
This paper explores the advancements and applications of language models in healthcare, focusing on their clinical use cases. It examines the evolution from early encoder-based systems requiring extensive fine-tuning to state-of-the-art large language and multimodal models capable of integrating text and visual data through in-context learning. The analysis emphasizes locally deployable models, which enhance data privacy and operational autonomy, and their applications in tasks such as text generation, classification, information extraction, and conversational systems. The paper also highlights a structured organization of tasks and a tiered ethical approach, providing a valuable resource for researchers and practitioners, while discussing key challenges related to ethics, evaluation, and implementation.
1 Introduction
@@
The advances in the area of artificial intelligence (AI) in recent years has opened countless opportunities for various sectors, including healthcare. The potential influence of AI is a subject of debate concerning its impact on humanity. Leading AI experts have called for caution, evidenced by an open letter urging a pause in the expansion of advanced AI models, which reflects growing concerns among policymakers and the public about the ethical, social, and economic ramifications of AI. While some argue that AI can bring substantial advances in efficiency and effectiveness across many sectors, others fear it could exacerbate inequalities, displace jobs, and challenge societal norms [1]. While AI in healthcare has an extensive history of research [2], the emergence of advanced foundational models [3] such as GPT family[4], Gemini [5], and a series of open models like the Llama [6], offers unprecedented perspectives for the transformation of the healthcare sector.
This review paper synthesizes and critically examines the landscape of language models in the medical domain, with a particular emphasis on their clinical applications. Compared to other reviews in this area, it highlights the following aspects:
  1. 1.
    A dedicated focus on clinical applications.
  2. 2.
    An emphasis on locally deployable models.
  3. 3.
    A structured organization based on NLP task categories, with appropriate justification for this framework.
  4. 4.
    A structured ethical assessment of reviewed tasks.
  5. 5.
    A detailed discussion of evaluation challenges.
Table 1 outlines how these distinguishing aspects compare to the focus areas of other reviews in the field.
{adjustwidth}
-2.25in0in Criterion Luo et al. (2024) [7] Meng et al. (2024) [8] Wang et al. (2023) [9] Thirunav- ukaras et al. (2023) [10] Yang et al. (2023) [11] Focus on clinical applications Partially Partially No Yes Yes Emphasis on locally deployable models Partially No Yes No Partially Structured by NLP task categories with justification Yes Partially Yes No Partially Structured ethical assessment of reviewed tasks No Partially No Partially Partially Detailed discussion of evaluation challenges Partially Partially Partially Partially Partially
Table 1: Unique aspects of this review compared to others in the field.
This review navigates the advancements from early encoder-based systems to state-of-the-art large language and multimodal models. It places a particular emphasis on locally deployable models, which enhance data privacy and adaptability within healthcare settings. Additionally, the structured exploration of ethical considerations and evaluation challenges provides a critical perspective on the current state of these models. By offering both a clear roadmap and valuable insights, this review serves as a comprehensive resource for newcomers and experienced interdisciplinary researchers, fostering innovation and promoting responsible implementation in this vital domain.
1.1 Definitions
Artificial Intelligence (AI): A multidisciplinary field aimed at creating systems capable of tasks that typically require human intelligence, such as problem-solving, learning, reasoning, and perception [12].
Machine Learning (ML): A branch of AI focused on developing algorithms and statistical models that enable computer systems to improve performance on specific tasks by learning from data and identifying patterns, rather than relying on explicit programming [12, 13].
Natural Language Processing (NLP): A branch of AI that deals with the interaction between computers and human language, allowing machines to understand, interpret, and generate human languages [12, 14].
Language Models: Probabilistic models that learn the statistical structure of word sequences, enabling the prediction or generation of text based on patterns observed in previously seen data [14].
Large Language Models (LLMs): Language models trained on extensive text corpora using deep learning architectures, enabling them to acquire broad linguistic knowledge and perform diverse natural language tasks [15, 14].
Generative Large Language Models (Generative LLMs): A subclass of LLMs that generate coherent and contextually relevant text by learning statistical patterns from large-scale language datasets, enabling tasks such as text generation, summarization, and dialogue creation.
Open LLMs: LLMs developed and distributed with open access to their architectures and weights, facilitating transparency, reproducibility, and community-driven improvements [16].
Foundation Models: Large-scale, general-purpose AI models, such as LLMs, trained on massive and diverse datasets. They serve as a base for a wide range of downstream tasks, often requiring minimal additional training or fine-tuning [3].
Embeddings: Dense vector representations of words, phrases, or other linguistic units that capture semantic relationships and similarity based on contextual usage in large corpora [14].
API (Application Programming Interface): In the context of LLMs, an API provides a set of tools and protocols that enable users to interact with a language model remotely, without requiring local deployment. Typically, API access abstracts away the model’s internals, offering only input-output functionality.
GPU (Graphics Processing Unit): A specialized processor originally designed for rendering graphics, now widely used to accelerate parallelizable computations, particularly in AI and ML tasks. Locally deployed models often require GPUs to run efficiently due to their ability to handle large-scale matrix operations.
@@ 2.1 Language Models
A critical milestone in the field of Natural Language Processing (NLP) was the introduction of the attention mechanism in neural machine translation [17]. This mechanism, linking the encoder and decoder in sequence-to-sequence models, paved the way for subsequent advancements by enabling the model to focus on different parts of the input sequence for each step of the output, substantially improving the handling of longer input sequences and complex dependencies. Notably, it led to the creation of the Transformer model [18], which exclusively relies on attention mechanisms. This innovation revolutionized not only the field of NLP but also the broader realms of AI and machine learning (ML).
The emergence of Transformer-based models has spurred the development of a wide array of language models, both in commercial and open formats. Among these, the GPT series [19] has advanced text generation capabilities while establishing the foundation for interactive applications. A notable milestone was the introduction of ChatGPT [20], a conversational agent that revolutionized chat-based language models by facilitating intuitive, human-like dialogue generation. The success of ChatGPT and similar systems [5, 21] underscored the adaptability of Transformer architectures and catalyzed the development of language models optimized for interactive and specialized applications, including those in the medical domain.
@@
Despite the simplicity of statistical approaches to language modeling and their initial lack of attention to the underlying rules of language, as critiqued by Chomsky [22], contemporary large language models (LLMs) like GPT-4 have demonstrated remarkable proficiency in a wide range of language understanding tasks [4], exhibiting emergent abilities and an advanced capacity to learn natural language patterns.
Modern language models have undergone a notable paradigm shift, moving from a pre-training and fine-tuning approach to embracing in-context learning (ICL) [23] and zero-shot learning [24]. Traditionally, language model development has followed a two-step process: pre-training and fine-tuning. During pre-training, a model learns general language representations from large text corpora in an unsupervised manner. For example, BERT [25] predicts masked tokens, while GPT-2 [19] predicts the next token in a sequence. The fine-tuning phase then adapts these pre-trained models to specific tasks (e.g., classification, summarization) using labeled data. Although effective, this approach often demands substantial task-specific datasets, which can limit its scalability for diverse downstream applications.
In contrast, zero-shot learning enables models to generalize to unseen tasks without requiring task-specific fine-tuning. Instead, the model leverages its pre-trained knowledge and interprets carefully designed prompts to perform tasks directly. For example, a zero-shot model can respond to a prompt such as “Summarize the key findings of this medical report” without being explicitly trained on clinical summarization datasets.
In-context learning enhances the capabilities of these models by enabling them to solve tasks using information provided within the input prompt. Unlike fine-tuning, ICL does not require updating the model’s weights. Instead, the model temporarily “learns” from examples in the prompt. For instance, the model can be provided with a series of medical cases that include patient symptoms, relevant background knowledge, and corresponding diagnoses. When presented with a new case, such as a patient with specific symptoms (e.g., coughing up phlegm and blood), the model can infer the most likely diagnosis based on the examples and background knowledge, without explicit task-specific fine-tuning [26].
Nevertheless, challenges remain, as highlighted by Mahowald et al. [27], who emphasize the gap between formal and functional linguistic competencies in LLMs. Formal linguistic competencies refer to a model’s ability to understand and generate syntactically and semantically correct language, enabling tasks like sentence completion, grammar correction, or summarization. In contrast, functional competencies involve applying language in practical, goal-oriented contexts, such as interpreting patient symptoms to suggest a diagnosis or deriving actionable steps from clinical guidelines. Achieving functional performance requires not only linguistic fluency but also reasoning and task-specific understanding, integrating broader contextual and intentional knowledge.
This gap can be analogized to human neuroscience, where distinct neural mechanisms underpin linguistic processing and reasoning [27]. Similarly, LLMs, trained predominantly on textual data, excel at linguistic mimicry but often lack the modularity and cognitive integration necessary for complex functional reasoning. To bridge this dissociation, Mahowald et al. propose integrating modular architectures or revising training processes to better align with these competencies.
For medical professionals, this gap underscores the limitations of LLMs in healthcare. While LLMs may summarize medical literature with linguistic precision, they may struggle to apply this information effectively in clinical scenarios, such as tailoring treatment plans based on patient-specific data. Recognizing these limitations is critical to ensuring LLMs are deployed as tools that complement, rather than replace, human expertise.
Notably, current research is shifting towards multimodal models that integrate multiple modalities, such as visual and textual data, into a single model [4, 5, 28], paving the way for more comprehensive AI solutions. While this review focuses on text-only language models, it briefly examines the ecosystem of multimodal models to highlight their emerging relevance. This exploration provides readers with starting points for further investigation into how these models expand the scope of AI applications, particularly in domains like medical imaging, where integrating textual and visual data is essential.
2.2 Locally-Deployable Models
Models can be categorized based on their data management and deployment strategies: locally-deployable models and API-based models. API-based models, such as GPT-4, require data transfer to third-party servers via web interfaces or APIs. In contrast, locally-deployable models run on an organization’s hardware, offering full data control and independence from external vendors.
In the medical domain, locally-deployable models provide critical advantages by ensuring sensitive medical data remains internal, complying with stringent data protection regulations, and enhancing operational autonomy by eliminating reliance on third-party providers.
Many locally-deployable models feature open weights and permissive usage terms [29]. This flexibility allows them to be customized for medical domain, optimized for specific types of inquiries, and seamlessly integrated with confidential datasets, while maintaining strict patient confidentiality.
@@ 2.2 Locally-Deployable Models
The primary limitations of locally-deployable models are their substantial hardware requirements and the technical expertise needed for deployment. Advanced models often comprise a significant number of parameters, typically represented as floating-point values. While a higher parameter count generally improves reasoning, pattern recognition, and linguistic capabilities [61], it also increases storage demands and necessitates more powerful GPUs for both inference and fine-tuning.
Recent advancements have addressed the computational challenges of large models through parameter-efficient fine-tuning and memory optimization. Techniques like Low-Rank Adaptation (LoRA) [62] update only low-rank parameters, reducing the need to adjust all model weights. Quantization [63] further minimizes memory usage and accelerates inference, enabling broader hardware compatibility. Combining these, Quantized Low-Rank Adaptation (QLoRA) [64] efficiently fine-tunes large models while significantly lowering memory and computational requirements. For example, a 1-billion-parameter model using 16-bit precision requires around 2GB of GPU memory, whereas quantization can halve this demand. Additional innovations, such as Virtual Memory Stitching (VMS) [65], optimize GPU memory allocation, reducing fragmentation and usage. Flash Attention [66] further enhances hardware efficiency, expanding the feasibility of deploying larger models on on-premise systems.
Table 2 summarizes widely used locally-deployable textual and multimodal models (predominantly English-language) categorized by the number of parameters and their architectures. The table distinguishes between three primary architectures: encoder-only, decoder-only, and encoder-decoder, each of which brings its own strengths and use cases.
Encoder-only models (e.g., BERT [25], DeBERTa [32]) excel in understanding and classification tasks, consistently achieving strong performance on comprehension-oriented benchmarks [67, 68, 69]. Decoder-only models (e.g., GPT-Neo [30], Falcon [46]) are better suited for generative tasks, such as text completion and narrative generation. Encoder-decoder models (e.g., BART [35], Flan-T5 [37]) strike a balance between comprehension and generation. While multimodal models can extend the capabilities of any of these architectures by incorporating multiple modalities (e.g., text and vision), we deliberately omit these details to maintain the review’s focus on text-based systems.
General Performance Guidelines
General evaluation of encoder-only language models typically focuses on tasks that assess deep language understanding without requiring generative capabilities. Benchmarks such as GLUE [70] and SuperGLUE [69], encompassing tasks like natural language inference, sentiment analysis, paraphrase detection, and coreference resolution, are widely adopted. Many of these benchmarks maintain leaderboards to track state-of-the-art performance [71]. Since encoder-only models often need fine-tuning for specific downstream tasks, it is advisable to prioritize evaluation on benchmarks that reflect the intended application.
The performance of decoder-only models is typically measured by their ability to generate coherent text, comprehend and reason about given inputs, and follow instructions. Prominent benchmarks include MMLU [72] and MMLU-Pro [73], while the technical reports of models like GPT-3 [24] provide a comprehensive list of metrics tailored for generative, decoder-only models.
Encoder-decoder models, designed for sequence-to-sequence tasks, are usually evaluated on benchmarks that measure their capacity to transform input sequences into output sequences. Examples include SQuAD [67] for question answering and XSum [74] for summarization. Readers interested in a deeper understanding of encoder-decoder architectures and their performance considerations are encouraged to review the T5 paper [39].
Evaluation of multimodal models often involves visual question answering (VQA) benchmarks. For instance, TextVQA [75] focuses on text reading in natural images, DocVQA [76] targets document understanding, and ChartQA [77] assesses chart comprehension. The Gemini technical report [5] offers a comprehensive list of metrics for vision-and-language multimodal models.
@@ 2.3 Domain-specific Language Models in Healthcare
Domain-specific language models are models tailored for a narrow field or topic, enhancing their ability to produce more accurate and relevant responses within that specific area. Generally, there are two common methods to create a domain-specific model: one approach is to pre-train a model using a set of domain-specific documents, such as medical papers; another is to take a generically trained model and fine-tune or adapt it to the target domain.
The medical field, rich in unstructured textual data from sources such as Electronic Health Records (EHR) and other medical documentation, presents an ideal research domain for language models to address a wide array of challenges. These applications range from tasks previously tackled by NLP techniques, such as clinical acronym disambiguation, to those unattainable a decade ago. For example, medical conversational agents that can assist both patients and healthcare professionals are now emerging. In response to these advancements, the research community has been actively developing language models specifically designed for medical applications. The evolution of these models, from pre-training and fine-tuning strategies to creating open models capable of zero-shot learning, highlights the growing synergy between artificial intelligence and healthcare [110, 116, 135, 117].
@@
The debut of BioGPT has marked another milestone in biomedical NLP. This decoder-based model, based on GPT-2 and trained on millions of PubMed abstracts, has demonstrated proficiency in specialized tasks, including text classification, and biomedical text generation [107].
@@
AlpaCare, a medical instruction-tuned LLM trained on more than fifty thousand instruction-response pairs generated artificially using GPT-4, represents a notable example from the current generation of medical LLMs that can be utilized in the medical domain without the need for fine-tuning on downstream tasks [110]. Clinical Camel [117], another medical LLM fine-tuned from the Llama-2 model, excels in various medical tasks, ranging from clinical note creation to medical triaging. Despite its limitations, such as the potential for generating misleading content and the need for continuous updates, it marked a remarkable advancement in medical LLMs. The issue of needing continuous updates was addressed by the ChatDoctor [116] model. Based on the Llama family and fine-tuned with real-world patient-doctor conversations, this model, as part of an integrated framework, can retrieve information from external sources, a crucial capability in the medical field, particularly for addressing emerging diseases.
@@
Recent advancements in the field, particularly the development of multimodal models that work with both text and images, have substantially impacted the medical domain as well. The LLaVA-Med model [123] can answer open-ended questions about biomedical images, while the CheXagent model [127], designed specifically for chest X-ray interpretation, exemplifies the successful transition from general medical models to specialist models capable of tackling problems in narrow medical fields.
@@
Table 3 provides a comprehensive, though not exhaustive, summary of existing medical language models. It highlights current trends in the development of medical models and the evolution of training and experimental datasets utilized. The experimental datasets listed in the table offer valuable references for readers interested in evaluating performance on specific tasks within the medical domain.
3 Medical Applications of Language Models